# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ IV-G Multihead Latent Attention (MLA)
Meng et al. [74] address the communication bottlenecks in modern LLMs by introducing Multi-head Latent Attention (MLA). This innovative mechanism uses low-rank matrices in the key-value layers to compress latent states, thereby reducing the KV cache size and speeding up inference. MLA employs an up-projection matrix to preserve expressiveness, effectively trading additional computation for reduced communication overhead. The authors also demonstrate that while existing models typically use Group Query Attention (GQA), any GQA can be represented with MLA—though not vice versa—highlighting MLA’s superior flexibility. To facilitate broader adoption, the paper introduces TransMLA, a post-training conversion method that transforms widely used GQA-based pre-trained models (e.g., LLaMA, Qwen, and Mixtral) into MLA-based models, allowing further training to boost expressiveness without incurring extra cache costs.
@@
IV-H Multimodal Rotary Position Embedding (M-RoPE)
@@
Wang et al. [63] propose the Multimodal Rotary Position Embedding (M-RoPE), which is an innovative extension of the traditional 1D rotary position embedding designed to encode positional information across diverse modalities effectively. By decomposing the original rotary embedding into three distinct components—temporal, height, and width—M-RoPE adapts to the unique spatial and temporal structures inherent in text, images, and videos. Textual data functions identically to conventional 1D-RoPE by using uniform position IDs, whereas images assign separate height and width identifiers while maintaining a constant temporal ID. In the case of videos, the temporal component increments with each frame, while the height and width follow the same schema as images. Additionally, when handling multiple modalities within a single input, M-RoPE sequentially increments the maximum position ID of the preceding modality to initialize the next, ensuring coherent integration across channels. This design enhances the model’s ability to capture and generalize positional information and reduces the numerical range of position IDs, thereby enabling efficient extrapolation to longer sequences during inference.
IV-I Multi-Token Prediction (MTP)
@@
DeepSeek-R1 is a next-generation language model architecture that significantly accelerates inference by employing Multi-Token Prediction (MTP) [42] to generate multiple tokens in parallel. The design incorporates cross-depth residual connections, enabling deeper MTP layers to leverage features from earlier ones. It also features adaptive prediction granularity that dynamically adjusts the number of tokens predicted based on the complexity of the input. Furthermore, DeepSeek-R1 employs depth-aware loss weighting using a sigmoid-based function to prioritize updates at mid-range depths, and it uses memory-efficient parameter sharing through depth-conditioned routing to reduce redundancy across transformer layers. Optimized speculative decoding, which relies on probabilistic agreement checking rather than exact matches, further speeds up inference. These architectural innovations result in substantial empirical improvements, with a 22% faster training convergence, a 1.5× increase in generation speed, and an 18% enhancement in long-form perplexity over its predecessor, DeepSeek-V3, demonstrating its superiority in training and inference efficiency.
IV-J Reinforcement Learning
@@
Reinforcement Learning (RL) for LLMs is a training approach where the model learns to improve its outputs based on feedback rather than solely relying on static labeled data. In this setting, the LLM acts as an agent that generates responses in reaction to prompts and then receives a reward signal—derived either from human feedback or automated metrics—that reflects the quality and relevance of its output.
IV-J1 Anchored Preference Optimization (APO)
@@
D’Oosterlinck et al. [90] explore advanced alignment techniques for large language models by leveraging contrastive learning and preference optimization. The authors identify that more contrastive preference pairs provide a stronger learning signal and that alignment objectives with greater control yield better performance. To capitalize on these insights, they introduce Contrastive Learning from AI Revisions (CLAIR)—a synthetic data generation method that produces minimally revised, highly contrastive preference pairs using a stronger model (e.g., GPT-4-turbo)—and Anchored Preference Optimization (APO), an RLHF approach that fine-tunes model output probabilities with nuanced control over winning and losing responses. Evaluated on MixEval-Hard, their best model, trained on 32K CLAIR preference pairs with APO, improves Llama-3-8B-Instruct’s performance by 7.65%, closing 45% of the gap with GPT-4-turbo, and demonstrates superior results compared to traditional discriminative reward models. Their approach also highlights the benefits of on-policy data and controlled optimization (using RMSProp with a learning rate of $2\times 10^{-7}$ over 18 epochs), emphasizing that the contrastiveness of preference pairs is a major driver of these improvements.
@@
IV-J2 Reinforcement Learning with Verifiable Rewards (RLVR)
Lambert et al. [40] propose Reinforcement Learning with Verifiable Rewards (RLVR) in the Tülu 3 model, which is a novel approach that redefines alignment for language models on tasks with verifiable outcomes, such as mathematical problem-solving and instruction following. RLVR builds on traditional RLHF by replacing the typical reward model with a verification function that provides a binary signal—rewarding the model only when its generated response is objectively correct. This simple yet effective strategy leverages answer matching or constraint verification to offer clear, targeted feedback, thereby streamlining the training process. RLVR improves performance on specialized benchmarks like GSM8K and maintains robustness across diverse tasks when integrated into a generalist training pipeline. By extending previous self-improvement and RL with execution feedback methods, RLVR demonstrates that reinforcing only verifiably correct outputs can lead to significant gains in model reliability and precision.
@@ IV-J2 Reinforcement Learning with Verifiable Rewards (RLVR)
IV-J3 Reinforcement Learning from Human Feedback (RLHF)
Reinforcement Learning from Human Feedback (RLHF) is an approach that leverages human judgments to shape the reward function used in training reinforcement learning agents. In the work by Christiano et al. [13], the reward model $r_{\theta}(x,y)$ is trained using human comparisons. For each prompt $x$, two responses are provided: one preferred $y_{w}$ and one less preferred $y_{l}$. The training objective ensures the model assigns a higher score to $y_{w}$ than to $y_{l}$. This is achieved by minimizing the loss function
@@ IV-J3 Reinforcement Learning from Human Feedback (RLHF)
In a subsequent extension of this methodology, Ouyang et al. [24] applied RLHF to fine-tune large language models. Here, the language model’s policy $\pi_{\phi}(y\mid x)$ is adapted using a reward model while constraining the updated policy to remain close to a baseline obtained through supervised fine-tuning, denoted as $\pi_{\text{SFT}}(y\mid x)$. The overall objective for updating the policy is given by
@@
Hu et al. [91] introduce REINFORCE++, an enhanced variant of the classical REINFORCE algorithm designed to streamline Reinforcement Learning from Human Feedback (RLHF) for large language models. By integrating optimization techniques inspired by Proximal Policy Optimization (PPO) while eliminating the need for a critical network, REINFORCE++ achieves a more straightforward training pipeline with improved stability and reduced computational overhead. Notably, the algorithm incorporates token-level KL penalties to mitigate divergence at a granular level. It employs a PPO-style clipping mechanism to ensure stable policy updates without the complexity of maintaining a separate value function. Empirical evaluations demonstrate that REINFORCE++ reduces training time by approximately 30% compared to PPO and holds, or even surpasses, the performance of more complex methods such as GRPO, particularly in tasks like mathematical reasoning. These results underscore the potential of REINFORCE++ to provide more efficient and stable RLHF training, paving the way for more cost-effective and robust alignment of language models with human preferences.
IV-J4 Reinforcement learning with execution feedback (RLEF)
Meta team [92] presented a novel reinforcement learning approach that integrates execution feedback for iterative code synthesis, addressing a critical limitation in current LLM-based agents which typically rely on independent sampling. The proposed method termed Reinforcement Learning with Execution Feedback (RLEF), trains models to optimize not only for first-attempt success but also for error correction through a cyclical process: the model generates code in response to an instruction, the code is executed and evaluated against public test cases, and the resulting feedback is used to guide subsequent corrections until either the public tests are passed or a predefined limit is reached. A final reward signal, derived from private test cases, is then used in a PPO framework to further refine performance. Experimental benchmarks on competitive programming tasks demonstrate that this method yields new state-of-the-art results for both small (8B parameter) and significant (70B parameter) models, significantly reducing the number of samples required by an order of magnitude. Notably, applying RLEF to Llama 3.1 70B improved its performance on CodeContests from 27.5% to 40.1% with up to three iterative corrections. It even outperformed state-of-the-art models like AlphaCodium that leverage GPT-4.
@@ IV-J4 Reinforcement learning with execution feedback (RLEF)
IV-J5 Self-Correction via Reinforcement Learning
SCoRe (Self-Correction via Reinforcement Learning) [93] is a novel multi-turn online RL approach that dramatically enhances the self-correction abilities of large language models using entirely self-generated data, addressing the shortcomings of traditional supervised fine-tuning methods that often suffer from distribution mismatches and ineffective correction behaviors. By employing a two-stage process—first using REINFORCE with a KL-divergence constraint to generate high-reward revisions while preserving the original response, and then lifting this restriction to allow full correction guided by a shaped reward that prioritizes changes from incorrect to correct answers—SCoRe enables models like Gemini 1.0 Pro and 1.5 Flash to significantly improve their performance, achieving a 15.6% boost on the MATH benchmark and a 9.1% improvement on HumanEval. Key insights from this approach include the critical role of on-policy sampling in multi-turn self-correction, the superior performance of REINFORCE over alternative methods like STaR, and further gains (an additional 10.5% improvement) when combined with inference-time scaling techniques, ultimately demonstrating that iterative, reinforcement-based self-correction allows models to learn from their mistakes and enhance complex reasoning capabilities.
IV-J6 Constrained Generative Policy Optimization (CGPO)
Meta team [94] propose Constrained Generative Policy Optimization (CGPO), a novel post-training paradigm aimed at addressing key limitations of Reinforcement Learning from Human Feedback (RLHF) in multi-task learning scenarios. Traditional RLHF methods such as PPO and DPO often encounter challenges with reward hacking and balancing multiple, sometimes conflicting, objectives, typically requiring extensive hyper-parameter tuning guided by human intuition. In contrast, CGPO leverages a Mixture of Judges (MoJ) approach that combines rule-based and LLM-based evaluators to enforce constraints in a principled manner, ensuring that the resulting policy reaches a pareto-optimal trade-off across a broad array of objectives. The methodology incorporates a DPO warm-up phase and then applies constrained policy optimization via stratification, effectively mitigating reward hacking while enhancing performance. Empirical results demonstrate that CGPO significantly outperforms standard RLHF algorithms, achieving improvements of 7.4% on AlpacaEval-2 for general chat tasks, 12.5% on Arena-Hard for STEM and reasoning tasks, and consistent gains in domains such as math and coding. Although the approach demands higher computational resources due to the involvement of multiple reward models and judges, it offers strong theoretical guarantees and a plug-and-play solution for aligning general-purpose LLMs across diverse applications.
@@ IV-J6 Constrained Generative Policy Optimization (CGPO)
IV-J7 Process Reward Models
Zhang et al. [4] investigated Process Reward Models (PRMs) as a method for supervising the intermediate reasoning steps in mathematical problem-solving with LLMs. The authors highlight the inherent challenges in developing effective PRMs, particularly regarding data annotation and evaluation. Their experiments reveal that traditional Monte Carlo (MC) estimation-based data synthesis, which uses a completion model to assess the correctness of each reasoning step, often underperforms compared to approaches that incorporate LLM-as-a-judge and human annotations. Specifically, the study identifies that MC estimation tends to produce inaccurate step verification and that conventional Best-of-N (BoN) evaluation strategies introduce biases by over-rewarding outcomes even when the underlying reasoning process is flawed. The authors propose a consensus filtering mechanism that integrates MC estimation and LLM-as-a-judge evaluations to mitigate these issues. This ensures that only steps with consistent labels are retained for training. This combined approach enhances the reliability and accuracy of step-level error identification and significantly improves overall model performance and data efficiency. The paper culminates in releasing state-of-the-art PRMs in 7B and 72B parameter configurations, demonstrating superior performance over existing open-source alternatives and offering practical guidelines for future advancements in process supervision models for mathematical reasoning.
Let’s consider a mathematical problem $p$ along with a proposed solution $s$. Two reward modeling strategies to evaluate the correctness and quality of $s$ are: Observational Reward Model (ORM) and Process Reward Model (PRM) [111].
@@ IV-J7 Process Reward Models
Cui et al. [112, 113] introduced PRIME, an open-source online reinforcement learning (RL) framework that leverages implicit Process Reward Modelling (PRM) to significantly enhance the reasoning abilities of LLMs. Unlike traditional approaches, PRIME directly learns a Q-function that assigns token-level rewards, enabling dynamic and online updates based solely on outcome improvements. The algorithm initializes both the policy model and the PRM with a supervised fine-tuning (SFT) baseline, and then generates rollouts from multiple prompts. These rollouts are scored using an implicit PRM in conjunction with an outcome verifier, with only those prompts yielding moderate success rates retained to stabilize training. PRIME computes binary outcome rewards and token-level process rewards, applies advantage estimation via RLOO, and ultimately updates the policy using a PPO loss function. Empirical results are impressive: for instance, the PRIME 7B model achieved a pass@1 rate of 26.7% on AIME 2024—substantially higher than the 3.3% from SFT and 9.3% from GPT-4o—along with an average improvement of 16.7% across key mathematical reasoning benchmarks. Additionally, the online updating of the PRM mitigates reward hacking and enhances reward accuracy while filtering out too-easy or too-hard prompts, further stabilizing the training process.
@@
Setlur et al. [114] present a novel approach to enhancing reasoning in large language models by leveraging process reward models (PRMs) that provide step-by-step feedback throughout a multi-step reasoning process. Unlike traditional outcome reward models (ORMs) that only offer input at the final step, this method defines an effective process reward as the improvement—or ”progress”—in the likelihood of eventually producing a correct response, measured before and after each reasoning step. Crucially, this progress is assessed under a distinct ”prover” policy, which can be a weaker LLM that helps guide exploration and error correction. The authors theoretically characterize the qualities of good prover models and empirically validate their approach by training process advantage verifiers (PAVs) to predict these step-level advantages. Their experimental results demonstrate that employing PAVs in test-time search increases accuracy by more than 8% and achieves 1.5 to 5 times better compute efficiency than standard ORMs. Moreover, reinforcement learning with dense rewards from PAVs yields a 5- to 6-fold gain in sample efficiency and over a 6% improvement in accuracy.
@@
Lightman et al. [25] investigate how to process supervision, which provides feedback on each intermediate reasoning step, can significantly improve the reliability of large language models on complex tasks compared to outcome supervision, which only evaluates the final result. By training models on the challenging MATH dataset, the study demonstrates that process-supervised models—augmented with active learning to optimize human feedback—achieve a 78% solve rate on a representative subset, outperforming outcome reward models (72.4%) and effectively teaching models to verify each step in their chain-of-thought. The methodology involves fine-tuning a base GPT-4 model to output delimited reasoning steps, collecting 800,000 human-labeled step-level feedback entries (PRM800K) across 75,000 solutions to 12,000 problems, and then using reinforcement learning with these fine-grained rewards to guide the model toward more accurate and reliable reasoning.
IV-J8 Critique-out-Loud Reward Models

### v2

@@ VI-G Multihead Latent Attention (MLA)
Meng et al. [meng2025transmla] address the communication bottlenecks in modern LLMs by introducing Multi-head Latent Attention (MLA). This innovative mechanism uses low-rank matrices in the key-value layers to compress latent states, thereby reducing the KV cache size and speeding up inference. MLA employs an up-projection matrix to preserve expressiveness, effectively trading additional computation for reduced communication overhead. The authors also demonstrate that while existing models typically use Group Query Attention (GQA), any GQA can be represented with MLA, though not vice versa, highlighting MLA’s superior flexibility. To facilitate broader adoption, the paper introduces TransMLA, a post-training conversion method that transforms widely used GQA-based pre-trained models (e.g., LLaMA, Qwen, and Mixtral) into MLA-based models, allowing further training to boost expressiveness without incurring extra cache costs.
@@
VI-H Multimodal Rotary Position Embedding (M-RoPE)
@@
Wang et al. [wang2024] propose the Multimodal Rotary Position Embedding (M-RoPE), which is an innovative extension of the traditional 1D rotary position embedding designed to encode positional information across diverse modalities effectively. By decomposing the original rotary embedding into three distinct components temporal, height, and width M-RoPE adapts to the unique spatial and temporal structures inherent in text, images, and videos. Textual data functions identically to conventional 1D-RoPE by using uniform position IDs, whereas images assign separate height and width identifiers while maintaining a constant temporal ID. In the case of videos, the temporal component increments with each frame, while the height and width follow the same schema as images. Additionally, when handling multiple modalities within a single input, M-RoPE sequentially increments the maximum position ID of the preceding modality to initialize the next, ensuring coherent integration across channels. This design enhances the model’s ability to capture and generalize positional information and reduces the numerical range of position IDs, thereby enabling efficient extrapolation to longer sequences during inference.
VI-I Multi-Token Prediction (MTP)
@@
DeepSeek-R1 is a next-generation language model architecture that significantly accelerates inference by employing Multi-Token Prediction (MTP) [deepseekai2025] to generate multiple tokens in parallel. The design incorporates cross-depth residual connections, enabling deeper MTP layers to leverage features from earlier ones. It also features adaptive prediction granularity that dynamically adjusts the number of tokens predicted based on the complexity of the input. Furthermore, DeepSeek-R1 employs depth-aware loss weighting using a sigmoid-based function to prioritize updates at mid-range depths, and it uses memory-efficient parameter sharing through depth-conditioned routing to reduce redundancy across transformer layers. Optimized speculative decoding, which relies on probabilistic agreement checking rather than exact matches, further speeds up inference. These architectural innovations result in substantial empirical improvements, with a 22% faster training convergence, a 1.5× increase in generation speed, and an 18% enhancement in long-form perplexity over its predecessor, DeepSeek-V3, demonstrating its superiority in training and inference efficiency.
VI-J Reinforcement Learning
@@
Reinforcement Learning (RL) [cheng2023impact] for LLMs is a training approach where the model learns to improve its outputs based on feedback rather than solely relying on static labeled data. In this setting, the LLM acts as an agent that generates responses in reaction to prompts and then receives a reward signal derived either from human feedback or automated metrics that reflects the quality and relevance of its output [lim2023reinforcement].
VI-J1 Anchored Preference Optimization (APO)
@@
D’Oosterlinck et al. [d2024anchored] explore advanced alignment techniques for large language models by leveraging contrastive learning and preference optimization. The authors identify that more contrastive preference pairs provide a stronger learning signal and that alignment objectives with greater control yield better performance. To capitalize on these insights, they introduce Contrastive Learning from AI Revisions (CLAIR)—a synthetic data generation method that produces minimally revised, highly contrastive preference pairs using a stronger model (e.g., GPT-4-turbo)—and Anchored Preference Optimization (APO), an RLHF approach that fine-tunes model output probabilities with nuanced control over winning and losing responses. Evaluated on MixEval-Hard, their best model, trained on 32K CLAIR preference pairs with APO, improves Llama-3-8B-Instruct’s performance by 7.65%, closing 45% of the gap with GPT-4-turbo, and demonstrates superior results compared to traditional discriminative reward models. Their approach also highlights the benefits of on-policy data and controlled optimization (using RMSProp with a learning rate of $2\times 10^{-7}$ over 18 epochs), emphasizing that the contrastiveness of preference pairs is a major driver of these improvements.
@@
VI-J2 Reinforcement Learning with Verifiable Rewards (RLVR)
Lambert et al. [lambert2025tulu3pushingfrontiers] propose Reinforcement Learning with Verifiable Rewards (RLVR) in the Tülu 3 model, which is a novel approach that redefines alignment for language models on tasks with verifiable outcomes, such as mathematical problem-solving and instruction following. RLVR builds on traditional RLHF by replacing the typical reward model with a verification function that provides a binary signal rewarding the model only when its generated response is objectively correct. This simple yet effective strategy leverages answer matching or constraint verification to offer clear, targeted feedback, thereby streamlining the training process. RLVR improves performance on specialized benchmarks like GSM8K and maintains robustness across diverse tasks when integrated into a generalist training pipeline. By extending previous self-improvement and RL with execution feedback methods, RLVR demonstrates that reinforcing only verifiably correct outputs can lead to significant gains in model reliability and precision.
@@ VI-J2 Reinforcement Learning with Verifiable Rewards (RLVR)
VI-J3 Reinforcement Learning from Human Feedback (RLHF)
Reinforcement Learning from Human Feedback (RLHF) is an approach that leverages human judgments to shape the reward function used in training reinforcement learning agents. In the work by Christiano et al. [christiano2017deep], the reward model $r_{\theta}(x,y)$ is trained using human comparisons. For each prompt $x$, two responses are provided: one preferred $y_{w}$ and one less preferred $y_{l}$. The training objective ensures the model assigns a higher score to $y_{w}$ than to $y_{l}$. This is achieved by minimizing the loss function
@@ VI-J3 Reinforcement Learning from Human Feedback (RLHF)
In a subsequent extension of this methodology, Ouyang et al. [ouyang2022training] applied RLHF to fine-tune large language models. Here, the language model’s policy $\pi_{\phi}(y\mid x)$ is adapted using a reward model while constraining the updated policy to remain close to a baseline obtained through supervised fine-tuning, denoted as $\pi_{\text{SFT}}(y\mid x)$. The overall objective for updating the policy is given by
@@
Hu et al. [hu2025reinforce++] introduce REINFORCE++, an enhanced variant of the classical REINFORCE algorithm designed to streamline Reinforcement Learning from Human Feedback (RLHF) for large language models. By integrating optimization techniques inspired by Proximal Policy Optimization (PPO) while eliminating the need for a critical network, REINFORCE++ achieves a more straightforward training pipeline with improved stability and reduced computational overhead. Notably, the algorithm incorporates token-level KL penalties to mitigate divergence at a granular level. It employs a PPO-style clipping mechanism to ensure stable policy updates without the complexity of maintaining a separate value function. Empirical evaluations demonstrate that REINFORCE++ reduces training time by approximately 30% compared to PPO and holds, or even surpasses, the performance of more complex methods such as GRPO, particularly in tasks like mathematical reasoning. These results underscore the potential of REINFORCE++ to provide more efficient and stable RLHF training, paving the way for more cost-effective and robust alignment of language models with human preferences.
VI-J4 Reinforcement learning with execution feedback (RLEF)
Meta team [gehring2024rlef] presented a novel reinforcement learning approach that integrates execution feedback for iterative code synthesis, addressing a critical limitation in current LLM-based agents which typically rely on independent sampling. The proposed method termed Reinforcement Learning with Execution Feedback (RLEF), trains models to optimize not only for first-attempt success but also for error correction through a cyclical process: the model generates code in response to an instruction, the code is executed and evaluated against public test cases, and the resulting feedback is used to guide subsequent corrections until either the public tests are passed or a predefined limit is reached. A final reward signal, derived from private test cases, is then used in a PPO framework to further refine performance. Experimental benchmarks on competitive programming tasks demonstrate that this method yields new state-of-the-art results for both small (8B parameter) and significant (70B parameter) models, significantly reducing the number of samples required by an order of magnitude. Notably, applying RLEF to Llama 3.1 70B improved its performance on CodeContests from 27.5% to 40.1% with up to three iterative corrections. It even outperformed state-of-the-art models like AlphaCodium that leverage GPT-4.
@@ VI-J4 Reinforcement learning with execution feedback (RLEF)
VI-J5 Self-Correction via Reinforcement Learning
SCoRe (Self-Correction via Reinforcement Learning) [kumar2024training] is a novel multi-turn online RL approach that dramatically enhances the self-correction abilities of large language models using entirely self-generated data, addressing the shortcomings of traditional supervised fine-tuning methods that often suffer from distribution mismatches and ineffective correction behaviors. By employing a two-stage process first using REINFORCE with a KL-divergence constraint to generate high-reward revisions while preserving the original response, and then lifting this restriction to allow full correction guided by a shaped reward that prioritizes changes from incorrect to correct answers SCoRe enables models like Gemini 1.0 Pro and 1.5 Flash to significantly improve their performance, achieving a 15.6% boost on the MATH benchmark and a 9.1% improvement on HumanEval. Key insights from this approach include the critical role of on-policy sampling in multi-turn self-correction, the superior performance of REINFORCE over alternative methods like STaR, and further gains (an additional 10.5% improvement) when combined with inference-time scaling techniques, ultimately demonstrating that iterative, reinforcement-based self-correction allows models to learn from their mistakes and enhance complex reasoning capabilities.
VI-J6 Constrained Generative Policy Optimization (CGPO)
Meta team [xu2024perfect] propose Constrained Generative Policy Optimization (CGPO), a novel post-training paradigm aimed at addressing key limitations of Reinforcement Learning from Human Feedback (RLHF) in multi-task learning scenarios. Traditional RLHF methods such as PPO and DPO often encounter challenges with reward hacking and balancing multiple, sometimes conflicting, objectives, typically requiring extensive hyper-parameter tuning guided by human intuition. In contrast, CGPO leverages a Mixture of Judges (MoJ) approach that combines rule-based and LLM-based evaluators to enforce constraints in a principled manner, ensuring that the resulting policy reaches a pareto-optimal trade-off across a broad array of objectives. The methodology incorporates a DPO warm-up phase and then applies constrained policy optimization via stratification, effectively mitigating reward hacking while enhancing performance. Empirical results demonstrate that CGPO significantly outperforms standard RLHF algorithms, achieving improvements of 7.4% on AlpacaEval-2 for general chat tasks, 12.5% on Arena-Hard for STEM and reasoning tasks, and consistent gains in domains such as math and coding. Although the approach demands higher computational resources due to the involvement of multiple reward models and judges, it offers strong theoretical guarantees and a plug-and-play solution for aligning general-purpose LLMs across diverse applications.
@@ VI-J6 Constrained Generative Policy Optimization (CGPO)
VI-J7 Process Reward Models
Zhang et al. [zhang2025lessons] investigated Process Reward Models (PRMs) as a method for supervising the intermediate reasoning steps in mathematical problem-solving with LLMs. The authors highlight the inherent challenges in developing effective PRMs, particularly regarding data annotation and evaluation. Their experiments reveal that traditional Monte Carlo (MC) estimation-based data synthesis, which uses a completion model to assess the correctness of each reasoning step, often underperforms compared to approaches that incorporate LLM-as-a-judge and human annotations. Specifically, the study identifies that MC estimation tends to produce inaccurate step verification and that conventional Best-of-N (BoN) evaluation strategies introduce biases by over-rewarding outcomes even when the underlying reasoning process is flawed. The authors propose a consensus filtering mechanism that integrates MC estimation and LLM-as-a-judge evaluations to mitigate these issues. This ensures that only steps with consistent labels are retained for training. This combined approach enhances the reliability and accuracy of step-level error identification and significantly improves overall model performance and data efficiency. The paper culminates in releasing state-of-the-art PRMs in 7B and 72B parameter configurations, demonstrating superior performance over existing open-source alternatives and offering practical guidelines for future advancements in process supervision models for mathematical reasoning.
Let’s consider a mathematical problem $p$ along with a proposed solution $s$. Two reward modeling strategies to evaluate the correctness and quality of $s$ are: Observational Reward Model (ORM) and Process Reward Model (PRM) [wang2023math].
@@ VI-J7 Process Reward Models
Cui et al. [cui2024process, yuan2024implicitprm] introduced PRIME, an open-source online reinforcement learning (RL) framework that leverages implicit Process Reward Modelling (PRM) to enhance the reasoning abilities of LLMs significantly. Unlike traditional approaches, PRIME directly learns a Q-function that assigns token-level rewards, enabling dynamic and online updates based solely on outcome improvements. The algorithm initializes both the policy model and the PRM with a supervised fine-tuning (SFT) baseline, and then generates rollouts from multiple prompts. These rollouts are scored using an implicit PRM in conjunction with an outcome verifier, with only those prompts yielding moderate success rates retained to stabilize training. PRIME computes binary outcome rewards and token-level process rewards, applies advantage estimation via RLOO, and ultimately updates the policy using a PPO loss function. Empirical results are impressive: for instance, the PRIME 7B model achieved a pass@1 rate of 26.7% on AIME 2024 substantially higher than the 3.3% from SFT and 9.3% from GPT-4o along with an average improvement of 16.7% across key mathematical reasoning benchmarks. Additionally, the online updating of the PRM mitigates reward hacking and enhances reward accuracy while filtering out too-easy or too-hard prompts, further stabilizing the training process.
@@
Setlur et al. [setlur2024rewarding] present a novel approach to enhancing reasoning in large language models by leveraging process reward models (PRMs) that provide step-by-step feedback throughout a multi-step reasoning process. Unlike traditional outcome reward models (ORMs) that only offer input at the final step, this method defines an effective process reward as the improvement or ”progress” in the likelihood of eventually producing a correct response, measured before and after each reasoning step. Crucially, this progress is assessed under a distinct ”prover” policy, which can be a weaker LLM that helps guide exploration and error correction. The authors theoretically characterize the qualities of good prover models and empirically validate their approach by training process advantage verifiers (PAVs) to predict these step-level advantages. Their experimental results demonstrate that employing PAVs in test-time search increases accuracy by more than 8% and achieves 1.5 to 5 times better compute efficiency than standard ORMs. Moreover, reinforcement learning with dense rewards from PAVs yields a 5- to 6-fold gain in sample efficiency and over a 6% improvement in accuracy.
@@
Lightman et al. [lightman2023let] investigate how to process supervision, which provides feedback on each intermediate reasoning step, can significantly improve the reliability of large language models on complex tasks compared to outcome supervision, which only evaluates the final result. By training models on the challenging MATH dataset, the study demonstrates that process-supervised models augmented with active learning to optimize human feedback achieve a 78% solve rate on a representative subset, outperforming outcome reward models (72.4%) and effectively teaching models to verify each step in their chain-of-thought. The methodology involves fine-tuning a base GPT-4 model to output delimited reasoning steps, collecting 800,000 human-labeled step-level feedback entries (PRM800K) across 75,000 solutions to 12,000 problems, and then using reinforcement learning with these fine-grained rewards to guide the model toward more accurate and reliable reasoning.
VI-J8 Critique-out-Loud Reward Models